Assume that a safe, risk-free asset provides a return of 25.00%. On the other hand, you could invest in a risky asset with expected returns of 52.00% and standard deviation of 13.50%. To reduce risk, you decide to hold both risk-free and risky assets in portfolio Y. What is the equation for the budget line relating mean portfolio returns on the vertical axis and standard deviation of the portfolio return on the horizontal axis?

Choose one:

A. 6.00+2𝜎𝑥


B. 35.00+2𝜎𝑥


C. 25.00+2𝜎𝑥


D. 25.00+3𝜎𝑥


E. 15.00+4𝜎𝑥



Relative to this budget line, what can we say about portfolio Z with given information (standard deviation, mean return) = (14.50, 9)?


Choose one:

A. We prefer portfolio Z.

B. We do not prefer portfolio Z.

C. Portfolio Z is on the budget line.

D. We're indifferent between portfolio Z and the original portfolio Y.




1
Expert's answer
2021-10-19T12:18:34-0400

Standard deviation"=" 13.50%

Risk free"=" 25%

Return"=" 52%


We use the formula

"E_p=r+(\frac{E_M-r}{\sigma_M})\sigma x" "=" 25.00+"(\frac{52-25}{13.50})\sigma x" =25.00"+" 2"\sigma x"



For portfolio Z using the same formula:

25.00"+" "(\frac{9-25}{14.50})\sigma x" "=25.00-1.10\sigma x"

So the answer is B. We do not prefer portfolio Z.